Sam is a guest on the TV show Math Jeopardy, and has just chosen the $300 question in the category "Quadratic Equations." The answer is "The solutions are x=3 and x=-2." What question could Sam ask that would win the $300? Is there more than one possible correct question? Explain.
Answer by checkley71(8403)

ONE QUESTION WOULD BE:
WHAT IS THE VALUE OF X IN THE EQUATION X^2-X-6=0
SOLUTION:
(X-3)(X+2)=0
X-3=0
X=3 ANSWER.
X+2=0
X=-2 ANSWER.